AP12-17A  (Ratio analysis of two companies)

You have obtained the financial statements of A-Tec and Bi-Sci, two new companies in the high-tech industry. Both companies have just completed their second full year of operations. You have acquired the following information for an analysis of the companies (amounts in thousands):

A-Tec

Bi-Sci

2020

2019

2020

2019

Cash

$10

$0

$25

$25

Accounts receivable

195

140

120

100

Inventory

130

100

110

100

Prepaid expenses

5

5

5

5

Capital assets(net)

350

300

230

160

Current liabilities

110

125

50

50

Lon-term debt

200

220

0

0

Share capital - Common shares

100

100

220

220

Retained earning

280

100

220

120

Sales (all credit sales)

1,900

1,300

1,250

1,200

Cost of goods sold

1,250

900

910

900

Interest expense

20

22

-

-

Taxes (30%)

77

56

64

56

Net income

180

130

150

130

Required

a.  Calculate the following ratios for the two companies for the two years. For 2019, assume the current year amount is equal to the average where required.

  1. Current ratio
  2. Accounts receivable turnover
  3. Inventory turnover
  4. Debt to equity
  5. Interest coverage
  6. Gross margin
  7. Profit margin
  8. Return on assets
  9. Return on equity

b.  Write a brief analysis of the two companies based on the information given and the ratios calculated. Be sure to discuss issues of short-term liquidity, activity, solvency, and profitability. Which company appears to be the better investment for the shareholder? Explain. Which company appears to be the better credit risk for the lender? Explain. Is there any other information you would like to have to complete your analysis?

Required

a.  Prepare a common-size analysis of the 2020 statement of income data for First Ltd. and Supreme Ltd.

b.  Calculate the return on assets and the return on shareholders' equity for both companies.
Answer #1

a). I.Current Ratio - It means whether the company has enough resources to meet its short term debts and obligations. The ideal current ratio is between 1.2 to 2, below 1 is means the company does not have enough resources to meet its short term obligations.

Hence Current Ratio =  Current Assets/Current Liabilties.

as per the Question :-

Current Assets = Cash + Accounts Receivables+Inventory+Prepaid Expenses

Current Liabilities are given directly, so now we will solve for each company for two years,

A-Tech

2019 :-   

Current Assets = 0+140+100+5 = $245

Current Liabilities = $125

Current Ratio = $245/$125 = 1.96

2020:-

Current Assets= 10+195+130+5 = $340

Current Liabilities = $110

Current Ratio = $340/$110 = 3.09

BI-Sci

2019:-

Current Assets = 25+100+100+5 = $230

Current Liabilities = $50

Current Ratio = $230/$50 = 4.6

2020:-

Current Assets = 25+120+110+5 = $260

Current Liabilities = $50

Current Ratio = $260/$50 = 5.2

II) Accounts Receivables Turnover :- It is tool used to measure how efficiently the company is able to issue credit to its customers and collect funds from them in a timely manner. it is also known as debtors turnover ratio.

Therefore Accounts receivables ratio = Net Credit sales / Average Accounts receivables

Net Credit sales = Sales on credit - Sales returns - Sales Allowances

Average Accounts Receivables = (Beginning A/R + Ending A/R)  / 2.

For 2019 we don't have opening A/R, but in the question it is said that whatever the Accounts receivables are provided in 2019 is itself Average A/R.

A Tech

2019:-

Accounts Receivables for 2019 = Average A/R

Net Credit Sales = $1300

Average Accounts Receivables = $140

Accounts Receivables turnover ratio = $1300/$140 = 9.28

That is Average Accounts Receivables was collected in 39.33 days. (365/9.28)

2020:-

Average Accounts receivables = (140+195)/2 = $167.5

Net Credit Sales = $1900

A/R Turnover Ratio = $1900/$167.5 = 11.34

i.e Average A/R was Collected in 34.83 days (365/11.34)

Bi Sci

2019:-

Net Credit Sales = $1200

Average A/r receivables = $100

A/R Turnover Ratio = $1200/$100 = 12

Average A/R was collected in 30.41 days

2020:-

Net Credit Sales = $1250

Average Accounts Receivables = (100+120) / 2 = $110

A/R turnover Ratio = $1250/$110 = 11.36

Average A/R was collected in 32.31 days

III). Inventory Turnover Ratio :- This tool measures how many times a company sold its total average inventory during the year.

For 2019 Inventory = Average Inventory

Inventory Turnover Ratio = COGS/Average Inventory

Average Inventory = (Beginning Inventory+Closing Inventory)/2

A-Tech

2019:-

COGS = $900

Average Inventory = $100

Inventory Turnover Ratio = $900/$100 = 9

2020:-

COGS = $1250

Average Inventory = (100+130)/2 = $115

Inventory Turnover ratio = $1250/$115 = 10.86

Bi-Sci

2019:-

COGS = $900

Average Inventory = $100

Inventory turnover ratio = $900/$100 = 9

2020:-

COGS = $910

Average Inventory = (100+110) / 2 = $105

Inventory Turnover Ratio = $910/$105 = 8.66

IV). Debt to Equity :- Total Liabilities / Total Shareholders Equity

as per question,

Total Liabilties = Current Liabilities + Long Term Liabilities

Shareholders Equity = Common Shares + Retained Earnings

A-Tech

2019:-

Total Liabilities = $125+$220 = $345

Shareholders Equity = $100+$100 = $200

Debt to Equity ratio = $345/$200 = 1.72

2020:-

Total Liabilties = $110+$200 = $310

Shareholders Equity = $100+$280 = $380

Debt to Equity ratio = $310/$380 = 0.82

Bi-Sci

2019:-

Total Liabilities = $50 (only Current Liabilities, since no Long term)

Shareholders Equity = $220+$120 = $340

Debt to Equity Ratio = $50/$340 = 0.15

2020:-

Total Liabilities = $ 50 (Only Current Liabilities)

Shareholders Equity = $220+$220 = $440

Debt to Equity Ratio = $50/$440 = 0.11

INTEREST Coverage Ratio: €811 / Toterest Expense Since, EBIT is not given, we can Calculate the Same by backward Calculation

Particulars A-Tech Bi-Sci
2019 2020 2019 2020
Current ratio 1.96 3.09 4.6 5.2
A/R Turnover ratio 9.28 11.34 12 11.36
Inventory turnover Ratio 9 10.86 9 8.66
Debt to Equity Ratio 1.73 0.82 0.15 0.11
Interest Coverage Ratio 9.45 13.85 - -
Gross Margin $400 $650 $300 $340

As per the ratios, Bi sci is more liquid that A tech and bases on Debt equity ratio A tech is good because it will be able to clear its debts but in 2020 it went down to less than 1 which is not a good sign.

a company should have both outside debts and shareholders equity which will be good.

it would be better to invest in Atech.